OpenStudy (anonymous):

Construct a basis of R4 that consists of the following vectors and some of the standard basis vectors e1,e2,e3,e4 in R4

OpenStudy (anonymous):

\[\left(\begin{matrix}1 \\ 2\\3\\4\end{matrix}\right)\left(\begin{matrix}1 \\ 4\\6\\8\end{matrix}\right)\]